Gary is reviewing his investment transactions for the year and determines he has net short-term capital losses of $8,100 and net long-term capital gains of $9,100.

Below is Gary's tax information:

Gary's tax bracket = 24%

Long-term Capital Gains tax rate = 18%

On the above situation, what are the taxes owed (or saved) in the current tax year for Gary ?

Enter your answer below as a whole number and without any $ signs.
Answer #1

Ans:

Short term capital loss can be set off against long term capital gain

Gary's Net Income from capital gain

Long term capital gain $9100
Less: Short term capital loss ($8100)
Net capital gain (long term) $1000

Tax on Long term capital gain= $1000 * 18% = $180

Tax saving on Short term capital loss = $8100 * 24% = $1944 ( Short term capital gains are taxed as ordinary income)

Tax saving on Long term capital gain = ($9100 * 18%) - $180 = $1548

Total tax saving = $1548 + $1944 = $3492

Therefore, in current year, Gary owed tax of 180 on long term capital gain and saved tax of 3492 (without $ sign)